A rectangular box has a volume of 750,000 cubic centimeters. It is
1 meter wide and 1.5 meters long. How tall is the box?

Answer by Alan3354(69443)

A rectangular box has a volume of 750,000 cubic centimeters. It is
1 meter wide and 1.5 meters long. How tall is the box?
---------
1 meter = 100 cm
1.5 meters = 150 cm
Volume = l*w*h = 750000 cc
750000 = 100*150*h = 1500h
h = 750000/1500 = 500 cm = 0.5 meters